if a sine wave has a rms voltage of 12 volts what will be its peak-to-peak voltage


1
Expert's answer
2021-07-06T08:22:18-0400

Since we know that "V_{rms}=\frac{1}{\sqrt{2}}V_{max}" and "V_{peak\,to\,peak}=2V_{max}" , we find the amplitud values for the peak-to-peak voltage as:


"V_{peak\,to\,peak}=(2\sqrt{2})V_{rms}=2\sqrt{2}(12\,V)=33.94\,V"


In conclusion, the peak-to-peak voltage of the sine wave is 33.94 V.


Reference:

  • Young, H. D., & Freedman, R. A. (2015). University Physics with Modern Physics and Mastering Physics (p. 1632). Academic Imports Sweden AB.